calculate the spacing between dislocations in a tilt boundary in fcc copper crystal , when the angle of tilt is 10° (burgers vector = 2.6 a°)
1
Expert's answer
2020-07-20T15:08:09-0400

As we know,


"\text{sin}\frac{\theta}{2}=\frac{b}{2D},\\\space\\
D=\frac{b}{2\text{sin}\frac{\theta}{2}}=\frac{2.6\cdot10^{-10}}{2\text{sin}\frac{10°}{2}}=1.49\cdot10^{-9}\text{ m}."

The dislocation is 1.49 nm.